OpenStudy (anonymous):

Show that e^xsinx = 0 has a solution in the interval [pi/2, 3pi/2] I just have to show that it excists, but how do I do that?

OpenStudy (kirbykirby):

\(e^x\sin x=0\) means either factor equals 0 .. think of when you solve \((x-a)(x-b)=0\) \(e^x=0\) or \(\sin x =0 \), but \(e^x\) is never 0, so can you find the solution for \(\sin x = 0\) in that interval ?

OpenStudy (anonymous):

i dont get it...

OpenStudy (kirbykirby):

Let's say you have a quadratic that you are solving for: \(x^2-4=0\) You can solve for x by factoring: (using a difference of squares): \((x-2)(x+2)=0\) Now that answer is 0 if: \((x-2)=0 \) or \((x+2)=0\) Now apply this same idea to your equation: \(e^x\sin x=0\\ (e^x)(\sin x)=0\) This is equal to 0 if \((e^x)=0 \) or \((\sin x) =0\) But \(e^x\) i the exponential function, and there is NO \(x\) that will render \(e^x=0\) true. (If you that the exponential function, the function never crosses the x-axis!, it has no root a.k.a. no zero) So the only way to make your equation 0 is if \(\sin x =0\) and do you know any value of \(x\) that makes \(\sin x = 0 \) ?

OpenStudy (kirbykirby):

If you look at the exponential function graphically*

OpenStudy (kirbykirby):